# MagicPose4D: Crafting Articulated Models with Appearance and Motion Control

### v1

@@ Abstract
With the success of 2D and 3D visual generative models, there is growing interest in generating 4D content. Existing methods primarily rely on text prompts to produce 4D content, but they often fall short of accurately defining complex or rare motions. To address this limitation, we propose MagicPose4D, a novel framework for refined control over both appearance and motion in 4D generation. Unlike traditional methods, MagicPose4D accepts monocular videos as motion prompts, enabling precise and customizable motion generation. MagicPose4D comprises two key modules: (i) Dual-Phase 4D Reconstruction Module which operates in two phases. The first phase focuses on capturing the model’s shape using accurate 2D supervision and less accurate but geometrically informative 3D pseudo-supervision without imposing skeleton constraints. The second phase refines the model using more accurate pseudo-3D supervision, obtained in the first phase and introduces kinematic chain-based skeleton constraints to ensure physical plausibility. Additionally, we propose a Global-local Chamfer loss that aligns the overall distribution of predicted mesh vertices with the supervision while maintaining part-level alignment without extra annotations. (ii) Cross-category Motion Transfer Module leverages the predictions from the 4D reconstruction module and uses a kinematic-chain-based skeleton to achieve cross-category motion transfer. It ensures smooth transitions between frames through dynamic rigidity, facilitating robust generalization without additional training. Through extensive experiments, we demonstrate that MagicPose4D significantly improves the accuracy and consistency of 4D content generation, outperforming existing methods in various benchmarks. Equal Contribution    Equal Advise
@@ 1 Introduction
However, challenges lie in facilitating users to freely and precisely specify the articulated motion of a non-rigid 3D object, and the generation faithfully reflecting the prompts, in terms of accurately capturing desired object appearance, geometry, and motion. The main issues with current methods are the following: (i) Temporal inconsistency in 3D geometry: Most current video generation models fail to ensure temporal consistency of 3D geometry. This involves additional complexity when objects are articulated because the 3D shape and movement must change mutually consistently during motion. For instance, while using the existing methods for 4D animal generation, we have observed unnatural and implausible configurations, such as the number of limbs of the object changing across different frames. (ii) Limited to common motions: Current video generation models perform well in generating simple, subtle, and common actions, such as walking or small motions like shaking, but they struggle to satisfactorily generate more complex motions, e.g., involving large or uncommon movements such as "King Kong dancing Hip-Hop" and "pig running like a rabbit" as shown in Fig.6. (iii) Text is inadequate for accurately describing the details of motion: As shown in Tab.4, most existing methods use text as prompts to describe the desired motion, with some efforts Bahmani et al. (2024) allowing trajectory (root body transformation) control. However, they are unable to specify detailed motions, e.g., by showing a real-world animal’s movement (say, via a monocular short video).
@@
Dual-Phase 4D Reconstruction Module: Given the complexity, especially with articulated objects, we divide this module into two phases. First, we use image descriptors (e.g., segments, flow) as 2D supervision. Additionally, we employ relatively less accurate but geometrically informative pseudo-3D models as geometric priors (using image-to-3D models) to provide 3D supervision for 3D reconstruction. This phase does not impose constraints emanating from the articulated nature of the object, e.g., captured in the underlying skeletal structure and plausible changes in it during motion, allowing each part to learn arbitrary rotations ($\mathbf{R}$) and translations ($\mathbf{t}$) while focusing on learning the model’s shape. In the second phase, we ensure the physical plausibility of the motion by enforcing kinematic chain-based skeletal movement constraints. Additionally, we propose a Global-local Chamfer loss, which ensures that the overall distribution of predicted mesh vertices aligns with the supervision while maintaining part-level alignment, without requiring additional annotations. Cross-category Motion Transfer Module: This module achieves cross-species motion transfer by mapping the skeleton of one species to another by establishing joint and limb correspondences through a kinematic-chain-based representation of the skeleton. Our motion transfer module is non-training-based. It helps improve generalization and prevents the poor performance seen in many existing methods when tested on data with significant gaps from the training set. Also, we leverage dynamic rigidity Zhang et al. (2024b) to guarantee smoothness between frames, unlike the existing approaches Song et al. (2021, 2023), which perform frame-independent pose transfer.
@@
Figure 1: Overview of MagicPose4D, which takes motion prompts (monocular video or dynamic mesh sequence) and appearance prompts (text or image) to control the 4D content generation.
@@ 2 Related Work & Motivation
4D/3D Reconstruction. Significant advancements in 4D/3D reconstruction include the development of specialized parametric models such as SCAPE Allen et al. (2003) and SMPL Loper et al. (2023a) for human bodies, MANO Romero et al. (2022) for hand movements, FLAME Li et al. (2017) and EMOCA Daněček et al. (2022) for facial expressions, and SMAL Zuffi et al. (2017) for quadruped animals. However, these methods require predefined parameters that include skeletons and skinning weights, limiting their generalization to uncommon species and out-of-distribution data. Recent neural implicit-based methods Palafox et al. (2021); Yang et al. (2022b); Giebenhain et al. (2023); Zhang et al. (2024b); Yang et al. (2021b) offer promising alternatives by jointly learning static 3D meshes and time-varying parameters without predefined templates. However, they often fail with monocular videos containing sparse views, limiting their practical applicability. To address this, MagicPose4D introduces an innovative Dual-Phase 4D Reconstruction module, which reduces the requirements from multi-view videos to single-view videos and achieves robust reconstruction.
@@ 3 MagicPose4D
MagicPose4D accepts two distinct types of input prompts: (i) appearance prompts and (ii) motion prompts. Consistent with recent methods Blattmann et al. (2023b); Tang et al. (2023), both images and textual descriptions can function as appearance prompts, delineating the desired object and its visual characteristics. In a departure from existing approaches, MagicPose4D enables users to specify precise motions and trajectories by providing a video/mesh sequence that represents the anticipated movement.
@@ 3.1 Terminology and Overview
Figure 2: Overview of Dual-Phase 4D Reconstruction Module.
@@ 3.2 Dual-Phase 4D Reconstruction from video
Reconstructing 4D models from short monocular videos is a challenging task as it requires jointly learning numerous parameters, resulting in an extensive optimization space. Erroneous predictions of any parameter can trigger cascading failures. To address these challenges, we propose a Dual-Phase 4D Reconstruction module, which employs differentiated supervision across two distinct phases and focuses on learning diverse representations.
@@ 3.2.1 Model Articulation
Skinning Weights $\mathbf{W}$ is designed to represent the probabilities that each vertex corresponds to $B$ semi-rigid parts. In the first phase, following Yang et al. (2021b, 2022b), the skinning weights are modeled by the mixture of $B$ Gaussian ellipsoids as: $\mathbf{W}_{n,b}=\mathcal{F}e^{-\frac{1}{2}(\mathbf{X}_{n}-\mathbf{C}_{b})^{T}\mathbf{Q}_{b}(\mathbf{X}_{n}-\mathbf{C}_{b})},$ where $\mathcal{F}$ is the factor of normalization and precision matrices $\mathbf{Q}_{b}=\mathbf{V}_{b}^{T}\mathbf{\Lambda}_{b}\mathbf{V}_{b}$. In each Gaussian ellipsoid, $\mathbf{C}\in\mathbb{R}^{B\times 3}$ denotes Gaussian centers, $\mathbf{V}\in\mathbb{R}^{B\times 3\times 3}$ defines the orientation and $\mathbf{\Lambda}\in\mathbb{R}^{B\times 3\times 3}$ denotes the diagonal scale matrix. $\mathbf{X}_{n}$ is the 3D location of vertex n. In the second phase, as shown in Fig.3, a skeletonization module, is leveraged to obtain a skeleton for the canonical mesh, and following Tong et al. (2012); Baran and Popović (2007), skinning weights $\mathbf{W}$ are obtained by a heat diffusion process. This approach guarantees a more natural assignment of skinning weights to the bones, enhancing the realism of the skeletal animations.
@@
$\begin{split}\mathbf{X}_{n}^{t}=\mathbf{\tau}_{0}^{t}(\sum_{b=1}^{B}\mathbf{W}_{n,b}\mathbf{\tau}_{b}^{t})\mathbf{X}_{n}^{0},\end{split}\vspace{-0pt}$ (1)
@@ 3.2.2 Supervision and Losses
As shown in Fig. 2, given the input from a monocular video, we utilize segmentation models Kirillov et al. (2023); Zhang et al. (2024a) and optical flow prediction models Teed and Deng (2020) to obtain 2D supervision. Furthermore, we employ an image-to-3D model to independently predict meshes for each frame, serving as pseudo-3D supervision. A question arises: "Why not directly use an image-to-3D model to independently predict 3D meshes for each frame?" As illustrated in Fig.7 (c), objects always exhibit self-occlusion in the video, under which circumstances image-to-3D models typically fail to produce accurate results. For instance, some frames might only depict a camel with two visible legs, resulting in a 3D mesh sequence that does not effectively capture the action information portrayed in the video. Moreover, as demonstrated in Video 3 of the supplementary materials, meshes generated directly from the image-to-3D module are independent of one another, thus lacking temporal continuity and smoothness. Although the initial pseudo-3D supervision may not be very accurate, it still provides valuable geometric priors that help address the information loss caused by insufficient perspectives of the object in the video.
In the first phase, we blend both 2D and 3D supervision to optimize the mesh shape and leverage a reconstruction loss, which consists of silhouette loss, optical flow loss, texture loss, perceptual loss, smooth, motion, and symmetric regularizations, and global-local chamfer (GLC) Loss. In the second phase, we only leverage GLC Loss and regularization terms without 2D losses. The details of the 2D loss functions and regularizations will be described in the Appendix.6.5.
@@
The second level involves computing the weighted chamfer distances between $B$ parts. First, we calculate skinning weights $\mathbf{W},\mathbf{\hat{W}}\in\mathbb{R}^{N\times B}$ for $\mathbf{S}$ and $\mathbf{\hat{S}}$ via the heat diffusion process Tong et al. (2012); Baran and Popović (2007). Then, we perform an argmax across the $B$ dimension to achieve a part-wise decomposition of the whole body into $K$ parts, where $K$ is less than or equal to $B$. In practice, $K$ often equals $B$, but in cases where $K$ is less than $B$, the loss computation simply omits the non-existent parts. Subsequently, we calculate the part-level Chamfer distances between the $K$ pairs from predicted mesh $\mathbf{P}_{k},k\in\{0,...,K-1\}$ and Pseudo mesh $\mathbf{\hat{P}}_{k}$, as shown in the following equation:
@@
Figure 3: Overview of Cross-Category Motion Transfer Module.
@@ 3.3 Cross-Category Motion Transfer
The input to the skeletonization module is a mesh, along with an optional skeleton template. When the reference object is a commonly recognized form, such as a quadruped or a human, we utilize the corresponding skeleton template and embed it into the mesh. Following the methodology described in Baran and Popović (2007), we construct a coarse discrete representation to locate the approximate position of the skeleton within the internal space of the mesh. This process involves embedding the skeleton into a graph derived from the character’s internal volume and determining an optimal solution using the A* algorithm across all possible matches. When a template of the object is not available, we employ skeleton extraction methods Bærentzen and Rotenberg (2021); Zhang et al. (2024b) to extract the canonical skeleton of the reference object. After learning the reference motion represented by the pre-frame angles and the bone scale of the skeleton, we can readily compute the global-local transformation of the target object using forward kinematics, as illustrated in Fig. 3(b). We extract the canonical skeleton for the target object by inputting the canonical skeleton of the reference object, adhering to the method previously described following Baran and Popović (2007). Subsequently, blend skinning is employed to generate the deformed meshes.
@@ 4 Experiments
This section primarily covers the following parts: (i) 4D reconstruction results and comparisons; (ii) motion transfer results and comparisons; and (iii) our framework v.s. 4D generation. Detailed descriptions of the benchmarks (Sec.6.2), implementation (Sec.6.3), and as shown in Tab.5, more than $60$ video results (Sec.6.1) are provided in the Appendix.
Figure 4: Appearance and Motion Controlled 4D Generation.
4.1 Qualitative and Quantitative Comparison
4D Reconstruction. As depicted in Fig.5, we first compare the results of MagicPose4D with LASR Yang et al. (2021a) and BANMo Yang et al. (2022a). These methods fall short of achieving ideal reconstruction, due to a lack of structural information about the objects, since they learn from only a single monocular video, which contains sparse views. NeRF-based methods such as BANMo Yang et al. (2022b) and MagicPony Wu et al. (2023) require particularly dense views from multiple cameras or long videos to achieve decent results. However, MagicPose4D achieves good reconstructions from only sparse views.
Furthermore, for a comprehensive quantitative comparison, we examine the performance of MagicPose4D against the best existing methods, LASR Yang et al. (2021a) and ViSER Yang et al. (2021c), which perform best (in terms of 2D keypoint transfer accuracy) among existing methods using a single monocular video as input. As shown in Tab.1, MagicPose4D consistently outperforms both LASR and ViSER for all animal subjects with a notable margin. When we run the author-provided codes for LASR, and ViSER on DAVIS, and PlanetZoo, we find that their results are highly variable across different runs. The results we report for these cases are therefore the mean accuracies we have obtained over multiple runs of all three methods.
Figure 6: 4D Generation Comparison of MagicPose4D with Animate124 Zhao et al. (2023), Motion Transfer Comparison of MagicPose4D with 3D-CoreNet Song et al. (2021) and X-DualNet Song et al. (2023). Videos are in Sec.6.1.
Cross-Category Motion Transfer. We present motion transfer results from MagicPose4D in Fig. 4. The target identities and reference motion sequences can either be humanoid or animal subjects. MagicPose4D is able to learn and retarget the motion from dynamic mesh prompts, e.g., Hip-Hop Dancing, or monocular video prompts, e.g., Hands Up.
We further compare the motion transfer ability for mesh generation quality with recent methods 3D-CoreNet Song et al. (2021) and X-DualNet Song et al. (2023) in Fig. 6 and Tab.2 (c). Both methods use a deep neural network to learn latent shape codes to retarget the motions and are trained/evaluated on SMPL Loper et al. (2023b)(humanoid) and SMAL Zuffi et al. (2017)(animal). Since they do not include disentangled components or shape deformation modules to represent the structural information of 3D meshes, these methods rely heavily on pre-processed training data with high-quality mesh annotations. Due to these constraints, these baselines cannot generalize well to in-the-wild target identities or reference motions. In contrast, MagicPose4D generates temporally consistent and smooth motions, while strictly preserving the identity and appearance of the target mesh. It is also worth noting that because previous methods focus only on pose transfer without considering the motion trajectory, the generated object always stays at the same position. This can be observed in side-by-side video comparisons in the supplementary materials Sec.6.1.

### v2

@@ Abstract
With the success of 2D and 3D visual generative models, there is growing interest in generating 4D content. Existing methods primarily rely on text prompts to produce 4D content, but they often fall short of accurately defining complex or rare motions. To address this limitation, we propose MagicPose4D, a novel framework for refined control over both appearance and motion in 4D generation. Unlike current 4D generation methods, MagicPose4D accepts monocular videos or mesh sequences as motion prompts, enabling precise and customizable motion control. MagicPose4D comprises two key modules: (i) Dual-Phase 4D Reconstruction Module which operates in two phases. The first phase focuses on capturing the model’s shape using accurate 2D supervision and less accurate but geometrically informative 3D pseudo-supervision without imposing skeleton constraints. The second phase extracts the 3D motion (skeleton poses) using more accurate pseudo-3D supervision, obtained in the first phase and introduces kinematic chain-based skeleton constraints to ensure physical plausibility. Additionally, we propose a Global-local Chamfer loss that aligns the overall distribution of predicted mesh vertices with the supervision while maintaining part-level alignment without extra annotations. (ii) Cross-category Motion Transfer Module leverages the extracted motion from the 4D reconstruction module and uses a kinematic-chain-based skeleton to achieve cross-category motion transfer. It ensures smooth transitions between frames through dynamic rigidity, facilitating robust generalization without additional training. Through extensive experiments, we demonstrate that MagicPose4D significantly improves the accuracy and consistency of 4D content generation, outperforming existing methods in various benchmarks. https://magicpose4d.github.io/ Equal Contribution
@@ 1 Introduction
However, challenges lie in facilitating users to freely and precisely specify the articulated motion of a non-rigid 3D object, and the generation faithfully reflecting the prompts, in terms of accurately capturing desired object appearance, geometry, and motion. The main issues with current methods are the following: (i) Temporal inconsistency in 3D geometry: Most current video generation models fail to ensure temporal consistency of 3D geometry. This involves additional complexity when objects are articulated because the 3D shape and movement must change mutually consistently during motion. For instance, while using the existing methods for 4D animal generation, we have observed unnatural and implausible configurations, such as the number of limbs of the object changing across different frames. (ii) Limited to common motions: Current video generation models perform well in generating simple, subtle, and common actions, such as walking or small motions like shaking, but they struggle to satisfactorily generate more complex motions, e.g., involving large or uncommon movements such as "King Kong dancing Hip-Hop" and "pig running like a rabbit" as shown in Fig.5. (iii) Text is inadequate for accurately describing the details of motion: As shown in Tab.4, most existing methods use text as prompts to describe the desired motion, with some efforts Bahmani et al. (2024a) allowing trajectory (root body transformation) control. However, they are unable to specify detailed motions, e.g., by showing a real-world animal’s movement (say, via a monocular short video).
@@
Dual-Phase 4D Reconstruction Module: Given the complexity, especially with articulated objects, we divide this module into two phases. First, we use image descriptors (e.g., segments, flow) as 2D supervision. Additionally, we employ relatively less accurate but geometrically informative pseudo-3D models as geometric priors (using image-to-3D models) to provide 3D supervision for 3D reconstruction. This phase does not impose constraints emanating from the articulated nature of the object, e.g., captured in the underlying skeletal structure and plausible changes in it during motion, allowing each part to learn arbitrary rotations ($\mathbf{R}$) and translations ($\mathbf{t}$) while focusing on learning the model’s shape. In the second phase, we ensure the physical plausibility of the motion by enforcing kinematic chain-based skeletal movement constraints. Additionally, we propose a Global-local Chamfer loss, which ensures that the overall distribution of predicted mesh vertices aligns with the supervision while maintaining part-level alignment, without requiring additional annotations. Cross-category Motion Transfer Module: This module achieves cross-species motion transfer by mapping the skeleton of one species to another by establishing joint and limb correspondences through a kinematic-chain-based representation of the skeleton. Our motion transfer module is non-training-based. It helps improve generalization and prevents the poor performance seen in many existing methods when tested on data with significant gaps from the training set. Also, we leverage dynamic rigidity Zhang et al. (2024c) to guarantee smoothness between frames, unlike the existing approaches Song et al. (2021; 2023), which perform frame-independent pose transfer.
@@
Figure 1: Overview of MagicPose4D. MagicPose4D takes motion prompts (monocular video or dynamic mesh sequence) and appearance prompts (text or image) to control 4D content generation. The dual-phase 4D reconstruction module extracts motion references, while the cross-category motion transfer module applies these motions to different target objects while ensuring temporal consistency.
@@ 2 Related Work & Motivation
4D/3D Reconstruction. Significant advancements in 4D/3D reconstruction include the development of specialized parametric models such as SCAPE Allen et al. (2003) and SMPL Loper et al. (2023a) for human bodies, MANO Romero et al. (2022) for hand movements, FLAME Li et al. (2017) and EMOCA Daněček et al. (2022) for facial expressions, and SMAL Zuffi et al. (2017) for quadruped animals. However, these methods require predefined parameters that include skeletons and skinning weights, limiting their generalization to uncommon species and out-of-distribution data. Recent neural implicit-based methods Palafox et al. (2021); Yang et al. (2022b); Giebenhain et al. (2023); Zhang et al. (2024c); Yang et al. (2021b) offer promising alternatives by jointly learning static 3D meshes and time-varying parameters without predefined templates. However, they often fail with monocular videos containing sparse views, limiting their practical applicability. To address this, MagicPose4D introduces an innovative Dual-Phase 4D Reconstruction module, which reduces the requirements from multi-view videos to single-view videos and achieves robust reconstruction.
@@ 3 MagicPose4D
MagicPose4D accepts two distinct types of input prompts: (i) appearance prompts and (ii) motion prompts. Consistent with recent methods Blattmann et al. (2023c); Tang et al. (2023), both images and textual descriptions can function as appearance prompts, delineating the desired object and its visual characteristics. In a departure from existing approaches, MagicPose4D enables users to specify precise motions and trajectories by providing a video/mesh sequence that represents the anticipated movement.
@@ 3.1 Terminology and Overview
Figure 2: Overview of Dual-Phase 4D Reconstruction Module.Given a monocular video, the Reconstruction Module extracts a physically plausible skeleton sequence by reconstructing a 3D mesh sequence. The first phase focuses on capturing the object’s shape, using a non-kinematic skeleton with learnable skinning weights for greater deformation flexibility. The second phase refines the skeleton for motion transfer, employing kinematic chains and heat diffusion-based skinning weights to ensure structural plausibility. Supervision shifts from a mix of 2D and pseudo-3D in the first phase to solely 3D pseudo-supervision in the second phase.
@@ 3.2 Dual-Phase 4D Reconstruction from video
Extracting the 3D motion of an articulated object directly from a monocular short video is extremely challenging. MagicPose4D addresses this issue by reconstructing the object and extracting its motion by learning its physically plausible skeleton.
Reconstructing 4D models from short monocular videos is also a challenging task as it requires jointly learning numerous parameters, resulting in an extensive optimization space. Erroneous predictions of any parameter can trigger cascading failures. To address these challenges, we propose a Dual-Phase 4D Reconstruction module, which employs differentiated supervision across two distinct phases and focuses on learning diverse representations.
@@ 3.2.1 Model Articulation
Skinning Weights $\mathbf{W}$ is designed to represent the probabilities that each vertex corresponds to $B$ semi-rigid parts. In the first phase, following Yang et al. (2021b; 2022b), the skinning weights are modeled by the mixture of $B$ Gaussian ellipsoids as: $\mathbf{W}_{n,b}=\mathcal{F}e^{-\frac{1}{2}(\mathbf{X}_{n}-\mathbf{C}_{b})^{T}\mathbf{Q}_{b}(\mathbf{X}_{n}-\mathbf{C}_{b})},$ where $\mathcal{F}$ is the factor of normalization and precision matrices $\mathbf{Q}_{b}=\mathbf{V}_{b}^{T}\mathbf{\Lambda}_{b}\mathbf{V}_{b}$. In each Gaussian ellipsoid, $\mathbf{C}\in\mathbb{R}^{B\times 3}$ denotes Gaussian centers, $\mathbf{V}\in\mathbb{R}^{B\times 3\times 3}$ defines the orientation and $\mathbf{\Lambda}\in\mathbb{R}^{B\times 3\times 3}$ denotes the diagonal scale matrix. $\mathbf{X}_{n}$ is the 3D location of vertex n. In the second phase, as shown in Fig.3, a skeletonization module, is leveraged to obtain a skeleton for the canonical mesh, and following Tong et al. (2012); Baran & Popović (2007), skinning weights $\mathbf{W}$ are obtained by a heat diffusion process. This approach guarantees a more natural assignment of skinning weights to the bones, enhancing the realism of the skeletal animations.
@@
$\begin{split}\mathbf{X}_{n}^{t}=\mathbf{\tau}_{0}^{t}(\sum_{b=1}^{B}\mathbf{W}_{n,b}\mathbf{\tau}_{b}^{t})\mathbf{X}_{n}^{0},\end{split}$ (1)
@@ 3.2.2 Supervision and Losses
As shown in Fig. 2, given the input from a monocular video, we utilize segmentation models Kirillov et al. (2023); Zhang et al. (2024a) and optical flow prediction models Teed & Deng (2020) to obtain 2D supervision. Furthermore, we employ an image-to-3D model to independently predict meshes for each frame, serving as pseudo-3D supervision. A question arises: "Why not directly use an image-to-3D model to independently predict 3D meshes for each frame?" As illustrated in Fig.6 (c), objects always exhibit self-occlusion in the video, under which circumstances image-to-3D models typically fail to produce accurate results. For instance, some frames might only depict a camel with two visible legs, resulting in a 3D mesh sequence that does not effectively capture the action information portrayed in the video. Moreover, as demonstrated in Video 3 of the supplementary materials, meshes generated directly from the image-to-3D module are independent of one another, thus lacking temporal continuity and smoothness. Although the initial pseudo-3D supervision may not be very accurate, it still provides valuable geometric priors that help address the information loss caused by insufficient perspectives of the object in the video.
In the first phase, we blend both 2D and 3D supervision to optimize the mesh shape and leverage a reconstruction loss, which consists of silhouette loss, optical flow loss, texture loss, perceptual loss, smooth, motion, and symmetric regularizations, and global-local chamfer (GLC) Loss. In the second phase, we only leverage GLC Loss and regularization terms without 2D losses. The details of the 2D loss functions and regularizations will be described in the Appendix.A.6.
@@
The second level involves computing the weighted chamfer distances between $B$ parts. First, we calculate skinning weights $\mathbf{W},\mathbf{\hat{W}}\in\mathbb{R}^{N\times B}$ for $\mathbf{S}$ and $\mathbf{\hat{S}}$ via the heat diffusion process Tong et al. (2012); Baran & Popović (2007). Then, we perform an argmax across the $B$ dimension to achieve a part-wise decomposition of the whole body into $K$ parts, where $K$ is less than or equal to $B$. In practice, $K$ often equals $B$, but in cases where $K$ is less than $B$, the loss computation simply omits the non-existent parts. Subsequently, we calculate the part-level Chamfer distances between the $K$ pairs from predicted mesh $\mathbf{P}_{k},k\in\{0,...,K-1\}$ and Pseudo mesh $\mathbf{\hat{P}}_{k}$, as shown in the following equation:
@@
Figure 3: Overview of Cross-Category Motion Transfer Module. In the second phase of 4D Reconstruction, the skeletal motion of reference meshes is extracted. The skeletonization module derives the canonical skeleton (at time = 0), which remains fixed along with skinning weights, allowing pose control via pre-frame angles and bone scales. A skeleton template is embedded using a graph-based approach if it is available; otherwise, skeleton extraction methods are applied. The learned motion parameters are then transferred to the target object using forward kinematics, and blend skinning is applied to generate the deformed meshes.
@@ 3.3 Cross-Category Motion Transfer
The input to the skeletonization module is a mesh, along with an optional skeleton template. When the reference object is a commonly recognized form, such as a quadruped or a human, we utilize the corresponding skeleton template and embed it into the mesh. Following the methodology described in Baran & Popović (2007), we construct a coarse discrete representation to locate the approximate position of the skeleton within the internal space of the mesh. This process involves embedding the skeleton into a graph derived from the character’s internal volume and determining an optimal solution using the A* algorithm across all possible matches. When a template of the object is not available, we employ skeleton extraction methods Bærentzen & Rotenberg (2021); Zhang et al. (2024c) to extract the canonical skeleton of the reference object. After learning the reference motion represented by the pre-frame angles and the bone scale of the skeleton, we can readily compute the global-local transformation of the target object using forward kinematics, as illustrated in Fig. 3(b). We extract the canonical skeleton for the target object by inputting the canonical skeleton of the reference object, adhering to the method previously described following Baran & Popović (2007). Subsequently, blend skinning is employed to generate the deformed meshes.
@@ 4 Experiments
This section primarily covers the experimental results and comparison of the following tasks: (i) 4D generation; (ii) 4D reconstruction; and (iii) Motion Transfer. Detailed descriptions of the benchmarks (Sec. A.3), and implementation (Sec. A.4) are in the Appendix. Additional visual results can be found on our anonymous webpage: https://magicpose4d.github.io/ and Google Drive: link.
Figure 4: Appearance and Motion Controlled 4D Generation. MagicPose4D can take either dynamic mesh sequences or monocular videos as motion prompts. These reference motions can be transferred to both humanoid and animal target identities.
Table 1: User study of MagicPose4D on 4D Generation compared to 4Dfy, DMT+Stag4D, and SVD+Stag4D. Criteria for judgment: (1) Appearance matches appearance prompts, (2) Motion matches text description, and (3) The overall generation quality. S4D is short for Stag4D. Prompts of experiments are shown in Tab.5.
Method Exp-1 Exp-2 Exp-3 Exp-4 Exp-5 Exp-6 Exp-7 Exp-8 Exp-9 Exp-10 Exp-11 Exp-12 Exp-13 Exp-14 Exp-15 Ave.
MP4D 5.04 4.97 5.17 5.14 5.01 5.05 4.81 5.01 5.16 5.00 5.02 5.02 4.89 4.94 4.64 4.99
4Dfy 3.07 2.83 3.98 2.75 3.55 2.82 3.27 3.03 3.23 3.25 3.00 2.89 3.16 2.97 2.75 3.09
DMT+S4D 2.61 2.44 1.78 2.08 2.13 2.62 2.63 2.84 2.37 2.73 2.60 2.41 2.64 2.67 2.37 2.52
SVD+S4D 3.25 2.86 3.40 3.29 3.08 3.26 3.14 3.31 3.35 3.31 3.22 3.02 3.03 3.31 2.94 3.19
Method Exp-16 Exp-17 Exp-18 Exp-19 Exp-20 Exp-21 Exp-22 Exp-23 Exp-24 Exp-25 Exp-26 Exp-27 Exp-28 Exp-29 Exp-30 Ave.
MP4D 5.25 5.17 4.86 4.92 5.33 5.14 4.99 5.17 5.37 4.97 4.97 5.39 5.20 4.91 5.15 5.08
4Dfy 3.63 3.25 3.36 3.12 3.56 2.95 2.95 3.29 2.25 2.34 2.90 2.68 3.32 2.73 2.49 3.00
DMT+S4D 2.65 2.40 2.34 2.42 2.11 3.01 2.35 2.47 1.86 2.22 2.49 1.97 2.60 2.19 2.32 2.40
SVD+S4D 3.69 3.05 3.15 3.20 3.08 2.85 3.97 3.19 2.59 3.67 2.49 3.32 2.07 2.43 3.31 3.12
4.1 4D Generation
Current 4D generation methods can be categorized into the following types based on how motion is controlled: (1) using text to describe the desired motion; (2) using video to specify the motion, and (3) lacking motion control or only allowing limited path control as shown in Tab.4.
The first category includes methods like SVD Blattmann et al. (2023b)+Stag4D Zeng et al. (2024) (which uses text to control SVD to generate videos with the desired motion and use Stag4D for reconstruction), 4Dfy Bahmani et al. (2024b), Animate124 Zhao et al. (2023), AYG Ling et al. (2023). The second category includes methods such as MagicPose4D and DMT Yatim et al. (2024)+Stag4D (where DMT enables video-level pose transfer). The third category comprises methods like DG4D Ren et al. (2023) and TC4D Bahmani et al. (2024a), which provide partial or no motion control.
As shown in Fig. 4, MagicPose4D takes either monocular videos or mesh sequences as motion prompts, enabling precise control over the motion of generated 4D articulated assets. Additional results are presented in the videos provided in the supplementary material. To highlight the advantages of MagicPose4D, we compared it with state-of-the-art (SOAT) methods that allow motion control from each category. However, because current SOAT methods mainly showcase simple, common, imprecise, small-scale, and short-duration motions ($1-2$ seconds), they fail to adequately assess a model’s ability to control motion. Therefore, we designed tests using more complex, uncommon, precise, large-scale, and long-duration motions (up to $20$ seconds), such as "run like a rabbit" (instead of just "run") and "dancing hip hop" ($300$ frames motion prompts) to better evaluate motion control.